What is the ground state energy for an electron that is confined to a potentialwell with a width of 0.2 micrometer um?
1
Expert's answer
2020-07-20T06:21:13-0400

 Potential well of width (L) = 0.2 × 10–6 m

For ground state n = 1,


Energy = n2h2 / 8mL2 .


So , Energy = (1 )2. (6.63*10-34)2 / 8 * ( 9.1 *10-31 ) * ( 10-6 )2 .


Energy = ( 6.63 )2 * 10-25 / 8 * ( 9.1)


Energy = 0.603 *10-25 j = 6.03 * 10-26 j . (Answer ) .


Energy = 6.03 *10 -26 / 1.6 * 10-19 = 3.76 * 10 -7 ev . ( ANSWER .) .